A chemistry student needs 65.0g of tetrahydrofuran for an experiment. By consulting the CRC Handbook of Chemistry and Physics, the student discovers that the density of tetrahydrofuran is 0.8889gxcm^-3. Calculate the volume of tetrahydrofuran the student should pour out.
Round your answer to significant digits
1
Expert's answer
2019-12-03T08:25:53-0500

V = m/p = 65/0.8889 = 73,1241 ≈ 73 cm3